Find an equation of the line containing the given pair of points (-1,-3)and (-8,-9). The equation of the line in the slope-intercept form is y=______.
Answer by John10(297)

Hint:
First, you need to find the slope
m = (y2 - y1)/(x2 - x1)
then you apply the slope-intercept form:
y - yo = m( x - xo) where (xo, yo) is one of the given points.